How do I find the secant 195?
Answer by stanbon(75887)

secant(195 degrees) = 1/cos(195 degrees) = -1.0353
----------------
secant(195) = secant(180+15) = -sec(15) = -1/cos(60-45)
---
= -1/(cos(60)cos(45)+sin(60)sin(45)) = -1/0.9659 = -1.053
=======================
Cheers,
Stan H.